THE BROADBAND ROTATIONAL SPECTRUM AND GEOMETRY OF N_{2}⋯ICF_{3}

\underline{N. R. WALKER}, D. HIRD, School of Chemistry, Bedson Building, Newcastle University, Newcastleupon-Tyne, NE1 7RU, U.K.; A. C. LEGON, School of Chemistry, University of Bristol, Bristol, BS8 1TS, U.K..

The rotational spectra of two isotopologues of N_{2}⋯ICF_{3} have been measured between 7 and 18.5 GHz by chirped-pulse Fourier transform microwave spectroscopy. The rotational constant, B_{0}, the centrifugal distortion constants, D_{J} and D_{JK} , and the nuclear quadrupole coupling constant of ^{127}I are precisely determined for isotopologues containing ^{14}N_{2} and ^{15}N_{2} respectively. The complex is a symmetric top and contains a linear arrangement of C, I, N(1) and N(2) atoms. The length of the halogen bond between N(1) and I will be presented. Experiments which exploit a Balle-Flygare FTMW spectrometer are currently in progress to determine the nuclear quadrupole coupling constants of the ^{14}N nuclei.